A 4.00 g sample of a mixture of CaO and BaO is placed in a 1.00 L flask containing CO2 gas at a pressure of 730 torr and a temperature of 25°C. The CO2 reacts with the CaO and BaO forming CaCO3 and BaCO3. When the reaction is complete, the pressure of the remaining CO2 is 150 torr.
Calculate the number of moles of CO2 that reacted.
Calculate the mass percentage of CaO in the mixture.
1
Expert's answer
2019-09-25T05:44:09-0400
"n CO2 = PV\/RT"

Before reaction

P=730torr= 97332.36 Pa

V = 1L= 0.001 m^3

T= 25 C= 298K

nCO2=0.039 mol

After reaction

P=150torr=20000 Pa

T and V same

nCO2 = 0.008 mol

CO2 which reacted its nCO2 before reaction minus nCO2 after reaction

nCO2 = 0.031 mol


"\omega CaO = mCaO/m.mix*100%"

Let be x it's mol of CO2 which reacted with CaO, and y it's mol of which reacted with BaO. We have system of equation

x+y = 0.031

x*MCaO + y*MBaO = 4

When we solved this system x = 0.008

Then m CaO = 0.008* 56 = 0.449 g


"\omega CaO = 11.2 ℅"